# ———————————————————————————————————–

### v2

@@ III.1 Quantum postprocessing with a coherent absorber
where $|\chi^{\rm ss}\rangle\in\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ is a purification of the system stationary state, i.e. $\rho^{\rm ss}={\rm Tr}_{\rm abs}(|\chi^{\rm ss}\rangle\langle\chi^{\rm ss}|)$. This implies that in the stationary regime the output is decoupled from system and absorber, and is in the ”vacuum” state $|0\rangle^{\otimes n}$. We briefly recall a few expressions related to the construction of $V$ that will be useful later on, and refer to Lemma 4.1 in Godley and Guţă (2023) for more details; for clarity, in the following we will use the labels $S,A,N$ to indicate system, absorber and noise unit. Let us consider a spectral representation of $\rho^{\rm ss}$ and the corresponding purification:
@@
where $V_{kl}$ and $K_{l}$ are the Kraus operators of $V$ and $U$ considered without ampliation. The following Lemma prescribes the structure of the blocks $V_{kl}$. We first define the “recovery” channel Petz (1986, 1988); Barnum and Knill (2002) with Kraus operators $K^{\prime}_{i}=\sqrt{\rho}K_{i}^{*}\sqrt{\rho^{-1}}$ and note that they satisfy the normalisation condition and the recovery channel has $\rho$ as invariant state.
@@
Later on we will require that the system and absorber transition operator $\tilde{\mathcal{T}}(\cdot)=\sum_{k}\tilde{K}^{*}_{k}\cdot\tilde{K}_{k}$ is primitive, in addition to the system’s transition operator $\mathcal{T}$ satisfying the same property. At the moment we are not able to establish what is the connection between the two properties. However we performed extensive numerical simulations with randomly chosen QMC dynamics and corresponding absorbers which indicate that “generically” with respect to the original dynamics, for every primitive $\mathcal{T}$ there exists a corresponding absorber such that $\tilde{\mathcal{T}}$ is primitive. In fact, a stronger statement seems to hold, which is that the spectral gap of $\tilde{\mathcal{T}}$ is smaller or equal to that of $\mathcal{T}$ and one can always choose the absorber such that the two are equal. For more details on the absorber theory we refer to the recent paper Tsang (2024).
@@
Returning to the parameter estimation setting where $U=U_{\theta}$ depends on the unknown parameter $\theta$, we note that one cannot implement a coherent absorber which precisely matches the system dynamics. Instead, one can implement the absorber for an approximate value $\theta_{0}$ of $\theta$ and try to estimate the offset $\theta-\theta_{0}$ by measuring the output. This setting is closely related to that of displaced null measurements discussed in section II. Indeed, the joint state of system, absorber and output is pure for all $\theta$ and at $\theta=\theta_{0}$ it is of the product form $|\chi^{\rm ss}_{\theta_{0}}\rangle\otimes|0\rangle^{\otimes n}$, assuming that system and absorber are initially in the stationary state. Therefore, repeated standard basis measurements on the output units constitute a null measurement (in conjunction with a final appropriate measurement on system and absorber).
@@
The exact procedure for determining $\theta_{0}$ will be described in section VII and follows the important displacement prescription outlined in section II. For the moment it suffices to say that $\theta_{0}$ will be informed by the outcome of a preliminary estimation stage involving simple (non-optimal) measurements on the output (without post-processing), and it will converge to $\theta$ in the limit of large $n$. While for $\theta_{0}=\theta$ the output state is the ”vacuum”, for $\theta\neq\theta_{0}$ the output could be seen as carrying a certain amount of ”excitations” which increases with the parameter mismatch $|\theta-\theta_{0}|$.
@@
In section IV we show how these ”excitations” can be given a precise meaning by fashioning the output Hilbert space into a Fock space carrying modes labelled by certain ”excitation patterns”. In section V we show that from this perspective, the output state converges to a joint coherent state of the excitation pattern modes whose displacement depends linearly on $\theta-\theta_{0}$. This will allow us to devise a simple ”pattern counting” algorithm for estimating $\theta$ in section VII.
IV Translationally invariant modes in the output
@@
In this section we introduce the concept of translationally invariant modes (TIMs) of a spin chain. We show that in the limit of large chain size, certain translationally invariant states acquire the characteristic Fock space structure, and that the corresponding creation and annihilation operators satisfy the bosonic commutation relations. This construction will then be used in analysing the stationary Markov output state in section V.
@@
Let $\mathcal{P}$ denote the ordered set of all patterns, where the order is the usual one when regarding patterns as integer numbers in binary representation. Let ${\bf n}:\mathcal{P}\to\mathbb{N}$ be pattern counts ${\bf n}=(n_{\alpha})_{\alpha\in\mathcal{P}}$ such that all but a finite number of counts are zero, and let ${\bf n}!:=\prod_{\alpha\in\mathcal{P}}n_{\alpha}!$ and $|{\bf n}|:=\sum_{\alpha}n_{\alpha}$ the total number of patterns.
@@
$|n_{1}=2;n\rangle$ ::=:= $\frac{1}{\sqrt{2}}\left(A_{1}^{*}(n)\right)^{2}|\Omega_{n}\rangle$
@@
$|n_{11}=1;n\rangle$ ::=:= $A_{11}^{*}(n)|\Omega_{n}\rangle$
@@
V Limit distribution of quadratures and number operators.
In this section we analyse the structure of the output state obtained by post-processing the output of a QMC with a coherent absorber, as described in section III.1. Motivated by the fact that the statistical uncertainty scales as $1/\sqrt{n}$ we choose the absorber parameter $\theta_{0}$ to be fixed and known, and write the system parameter as $\theta=\theta_{0}+u/\sqrt{n}$, where $u$ is an unknown local parameter. For a more in depth motivation we refer to local asymptotic normality theory in appendix A.
Since the output state becomes stationary for long times, it is natural to focus on the state of the translationally invariant modes introduced in section IV. In Theorem 2 we show that asymptotically with $n$, the restricted state of these modes is a joint coherent state whose amplitude depends linearly on $u$, i.e. a Gaussian shift model. Moreover, Corollary 2 shows that the QFI of this model is equal to the QFI rate $f_{\theta_{0}}$ of the output state characterised in Theorem 1. This means that the TIMs capture the entire QFI of the output. Together with Theorem 3 of section VI, these results will be the theoretical underpinning the estimator proposed in section VII.
We consider the system and absorber together as an open system with space $\mathbb{C}^{D}=\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ with $D=d^{2}$, interacting with the noise units. The corresponding unitary is $W_{\theta}=V_{\theta_{0}}U_{\theta}$ where the absorber parameter is a fixed value $\theta_{0}$ (which later will be determined based on a preliminary estimation procedure) and $\theta$ is the unknown parameter to be estimated. We distinguish between the system Kraus operators $K_{\theta,i}=\langle i|U_{\theta}|0\rangle\in M(\mathbb{C}^{d})$ and the system and absorber Kraus operators $\tilde{K}_{\theta,i}=\langle i|W_{\theta}|0\rangle\in M(\mathbb{C}^{D})$, and similarly between the system transition operator $\mathcal{T}_{\theta}$ and the system and absorber transition operator $\tilde{\mathcal{T}}_{\theta}(X)=\sum_{i=0}^{1}\tilde{K}_{\theta,i}^{*}X\tilde{K}_{\theta,i}$.
@@
    $\tilde{\mathcal{T}}_{\theta}$ has a unique faithful invariant state $\tilde{\rho}^{\rm ss}_{\theta}$ and is aperiodic for $\theta$ in a neighborhood of $\theta_{0}$;
@@
Note that properties $1$ and $2$ in Hyothesis 2 imply that $\dot{\tilde{\rho}}^{\rm ss}=\tilde{\mathcal{R}}_{*}\dot{\tilde{\mathcal{T}}}_{*}(\tilde{\rho}^{\rm ss})$ where $\tilde{\mathcal{R}}$ is the Moore-Penrose inverse of ${\rm Id}-\tilde{\mathcal{T}}$. In addition, $\tilde{K}_{0}{|{\chi^{\rm ss}}\rangle}=\tilde{K}^{*}_{0}{|{\chi^{\rm ss}}\rangle}={|{\chi^{\rm ss}}\rangle}$. Indeed by construction $\tilde{K}_{0}|\chi^{\rm ss}\rangle=|\chi^{\rm ss}\rangle$ and $\tilde{K}_{1}|\chi^{\rm ss}\rangle=0$, and by applying $\tilde{K}_{0}^{*}\tilde{K}_{0}+\tilde{K}_{1}^{*}\tilde{K}_{1}=\mathbf{1}$ to $|\chi^{\rm ss}\rangle$ we get $\tilde{K}_{0}^{*}|\chi^{\rm ss}\rangle=|\chi^{\rm ss}\rangle$.
@@
To formulate our results below we need to introduce several superoperators acting on the system and absorber space. For $x\in M(\mathbb{C}^{D})$, we define
${\cal A}_{i}(x)=\begin{cases}\tilde{\mathcal{T}}(x)&i=0\\ \tilde{K}^{*}_{1}x\tilde{K}_{0}&i=1\\ \tilde{K}^{*}_{0}x\tilde{K}_{1}&i=-1,\\ \end{cases}$ (15)
@@
The upshot of this section is that (asymptotically in $n$) the statistical information of the output state is concentrated in the TIMs and the state’s restriction to the TIM Bosonic algebra is a coherent state. Formally, in order to optimally estimate the parameter, one would only need to measure the appropriate quadrature of the Gaussian shift model, as explained in section II. However, it is not obvious how to perform such a measurement, and the theoretical insight does not seem to help on the practical side. Surprisingly, it turns out that the standard sequential counting measurement is an effective joint measurement of all the TIMs’ number operators! This will be the main result of the next section, which in conjunction with the displaced null strategy discussed in section A, provides the ingredients of a counting-based estimation strategy.
VI Limit theorem for counting trajectories
@@
Let us fix any real number $10<\gamma<$ 1. Let ${\bf m}=\{m_{\alpha}\}_{\alpha\in\mathcal{P}}$ be a set of pattern excitation counts where all occupation numbers are zero except $(m_{\alpha^{(1)}},\dots,m_{\alpha^{(k)}})$, and define $B_{\bf m}(n)$ as the empty set for every $n$ strictly smaller than $\sum_{i=1}^{k}m_{\alpha^{(i)}}|\alpha^{(i)}|+(k-1)n^{\gamma}$ ($|\alpha|$ is the length of the pattern $\alpha$), otherwise $B_{\bf m}(n)$ is the set of all binary sequences obtained concatenating $m_{\alpha^{(1)}}$ copies of $\alpha^{(1)}$s, up to $m_{\alpha^{(k)}}$ copies of $\alpha^{(k)}$s in any order, and inserting 00s before, between and after them in order to reach a total length of $n$ and making sure that between two consecutive patterns there are at least $n^{\gamma}$ 00s.
The following Theorem shows that the distribution of the pattern counts ${\bf m}$ converges to a product of Poisson distributions with the same intensities as those of the number operators of the TIM in Theorem 2. This means that performing a standard output measurement and extracting the pattern counts provides an effective joint measurement of the number operators of the TIMs. This finding is essential in constructing an optimal estimator in section VII.
@@
For every positive constant $C>0$ and finite collection of excitation patterns counts ${\bf m}=(m_{\alpha^{(1)}},\dots,m_{\alpha^{(k)}})$, the following limit holds true:
$\lim_{n\rightarrow+\infty}\sup_{|u|<C}\left|\nu_{u,n}\left(B_{\bf m}(n)\right)-e^{-\lambda_{tot}u^{2}}\prod_{i=1}^{k}\frac{\left(\lambda_{\alpha^{(i)}}u^{2}\right)^{m_{\alpha^{(i)}}}}{m_{\alpha^{(i)}}!}\right|=0.$ (20)
@@
$\lambda_{tot}$ ::=:= $-\frac{1}{2}{\langle{\chi^{\rm ss}}|}(2\dot{{\cal B}}_{0*}{\cal R}_{0*}\dot{{\cal B}}_{0*}+\ddot{{\cal B}}_{0*})({|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|}){|{\chi^{\rm ss}}\rangle}$ (21)
@@
The previous result has some relevant consequences. Let us define the “pattern extraction” function which associates to each trajectory $\omega\in\{0,1\}^{n}$ a set of pattern counts $\{N_{\alpha,n}(\omega):\alpha\in\mathcal{P}\}\in\mathbb{N}^{\mathcal{P}}$, which is uniquely determined by the condition that $\omega$ is a maximal union of contiguous patterns separated by sequences of 00s of length at least $n^{\gamma}$ with a fixed $10<\gamma<$ 1; moreover, let us consider the stochastic process given by the infinite collection of independent random variables $\{N_{\alpha}:\alpha\in\mathcal{P}\}$ where $N_{\alpha}$ is a Poisson random variable with parameter $\lambda_{\alpha}u^{2}$.
@@
However, in the following we will be interested in local parameters with growing size, i.e. $|u|\leq n^{\epsilon^{\prime}}$ for some $20<\epsilon^{\prime}<1/$ 2. In this case we can show the following result.
@@
For $60<\epsilon^{\prime}<1/$ 6 and for every finite collection of excitation patterns counts ${\bf m}$ the following holds true:
$\lim_{n\rightarrow+\infty}\sup_{|u|\leq n^{\epsilon^{\prime}}}\left|\frac{\nu_{u,n}\left(B_{\mathbf{m}}(n)\right)}{e^{-\lambda_{tot}u^{2}}\prod_{i=1}^{k}\frac{\lambda_{\alpha^{(i)}}^{m_{i}}u^{2m_{i}}}{m_{i}!}}-1\right|=0.$
The Proof of Proposition 2 can be found in Appendix E as well. Upgrading this result to a weak convergence one similar to Theorem 3 remains the subject of future research.
VII Pattern counting estimator
In this section we describe our adaptive estimation scheme which exploits the asymptotic results presented in sections V and VI. The scheme involves four key ingredients:

### v3

@@ 3.1 Quantum postprocessing with a coherent absorber
where $|\chi^{\rm ss}\rangle\in\mathcal{H}_{s}\otimes\mathcal{H}_{a}$ is a purification of the system stationary state, i.e. $\rho^{\rm ss}={\rm Tr}_{\rm abs}(|\chi^{\rm ss}\rangle\langle\chi^{\rm ss}|)$. This implies that in the stationary regime the output is decoupled from system and absorber, and is in the ”vacuum” state $|0\rangle^{\otimes n}$. We briefly recall a few expressions related to the construction of $V$ that will be useful later on, and refer to Lemma 4.1 in [115] for more details; for clarity, in the following we will use the labels $S,A,N$ to indicate system, absorber and noise unit. Let us consider a spectral representation of $\rho^{\rm ss}$ and the corresponding purification:
@@
where $\mathbf{1}_{S}\otimes V_{kl}:={\langle{k_{N}}|}V{|{l_{N}}\rangle}$ and $K_{l}\otimes\mathbf{1}_{A}:={\langle{l_{N}}|}U{|{0_{N}}\rangle}$ are the Kraus operators of $V$ and $U$, respectively. The following Lemma prescribes the structure of the blocks $V_{kl}$. We first define the “recovery” channel [133, 134, 135] with Kraus operators $K^{\prime}_{i}=\sqrt{\rho}K_{i}^{*}\sqrt{\rho^{-1}}$ and note that they satisfy the normalisation condition and the recovery channel has $\rho$ as invariant state.
@@
Later on we will require that the system and absorber transition operator $\tilde{\mathcal{T}}(\cdot)=\sum_{k}\tilde{K}^{*}_{k}\cdot\tilde{K}_{k}$ is primitive, in addition to the system’s transition operator $\mathcal{T}$ satisfying the same property. At the moment we are not able to establish what is the connection between the two properties. However we performed extensive numerical simulations with randomly chosen QMC dynamics and corresponding absorbers which indicate that “generically” with respect to the original dynamics, for every primitive $\mathcal{T}$ there exists a corresponding absorber such that $\tilde{\mathcal{T}}$ is primitive. In fact, a stronger statement seems to hold, which is that the spectral gap of $\tilde{\mathcal{T}}$ is smaller or equal to that of $\mathcal{T}$ and one can always choose the absorber such that the two are equal. For more details on the absorber theory we refer to the recent paper [136].
@@
Returning to the parameter estimation setting where $U=U_{\theta}$ depends on the unknown parameter $\theta$, we note that one cannot implement a coherent absorber which precisely matches the system dynamics. Instead, one can implement the absorber for an approximate value $\theta_{0}$ of $\theta$ and try to estimate the offset $\theta-\theta_{0}$ by measuring the output. This setting is closely related to that of displaced null measurements discussed in section 2. Indeed, the joint state of system, absorber and output is pure for all $\theta$ and at $\theta=\theta_{0}$ it is of the product form $|\chi^{\rm ss}_{\theta_{0}}\rangle\otimes|0\rangle^{\otimes n}$, assuming that system and absorber are initially in the stationary state. Therefore, repeated standard basis measurements on the output units constitute a null measurement (in conjunction with a final appropriate measurement on system and absorber).
Once again, let us consider the discretisation of the simple continuous time example in equation (6); using Lemma 1, one obtains the following Kraus operators for the system and the absorber together (the first factor corresponds to the system and the second one to the absorber):
$\widetilde{K}_{0}=e^{-i\omega_{t}\sigma_{z}}\otimes e^{i\omega_{t}\sigma_{z}}(\gamma_{t}\mathbf{1}+(1-\gamma_{t})\sigma_{x}\otimes\sigma_{x}),$
$\widetilde{K}_{1}=\sqrt{\gamma_{t}(1-\gamma_{t})}e^{-i\omega_{t}\sigma_{z}}\otimes W(\mathbf{1}-\sigma_{x}\otimes\sigma_{x}),$
where $W$ is an arbitrary unitary operator. One can check that the purification ${|{\chi^{\rm ss}}\rangle}=\frac{1}{\sqrt{2}}({|{00}\rangle}+{|{11}\rangle})$ of the unique invariant state for the system satisfies $\widetilde{K}_{0}{|{\chi^{\rm ss}}\rangle}={|{\chi^{\rm ss}}\rangle}$ and $\widetilde{K}_{1}{|{\chi^{\rm ss}}\rangle}=0$ (which implies that it is an invariant state for the dynamics of system and absorber together); In the continuous time counterpart, this would correspond to the following Hamiltonian $\widetilde{H}$ and jump operator $\widetilde{L}$ for the system and absorber dynamics:
$\widetilde{H}=\omega(\sigma_{z}\otimes\mathbf{1}-\mathbf{1}\otimes\sigma_{z}),\quad\widetilde{L}=\sqrt{\gamma}\mathbf{1}\otimes W(\mathbf{1}-\sigma_{x}\otimes\sigma_{x}).$
By choosing $W=-\sigma_{x}$ we recover the absorber parameters prescribed in [116] and the system-absorber dynamics can be described as cascaded input-output dynamics.
@@
The exact procedure for determining $\theta_{0}$ will be described in section 7 and follows the important displacement prescription outlined in section 2. For the moment it suffices to say that $\theta_{0}$ will be informed by the outcome of a preliminary estimation stage involving simple (non-optimal) measurements on the output (without post-processing), and it will converge to $\theta$ in the limit of large $n$. While for $\theta_{0}=\theta$ the output state is the ”vacuum”, for $\theta\neq\theta_{0}$ the output could be seen as carrying a certain amount of ”excitations” which increases with the parameter mismatch $|\theta-\theta_{0}|$.
@@
In section 4 we show how these ”excitations” can be given a precise meaning by fashioning the output Hilbert space into a Fock space carrying modes labelled by certain ”excitation patterns”. In section 5 we show that from this perspective, the output state converges to a joint coherent state of the excitation pattern modes whose displacement depends linearly on $\theta-\theta_{0}$. This will allow us to devise a simple ”pattern counting” algorithm for estimating $\theta$ in section 7.
4 Translationally invariant modes in the output
@@
In this section we introduce the concept of translationally invariant modes (TIMs) of a spin chain. We show that in the limit of large chain size, certain translationally invariant states acquire the characteristic Fock space structure, and that the corresponding creation and annihilation operators satisfy the bosonic commutation relations. This construction will then be used in analysing the stationary Markov output state in section 5.
@@
Let $\mathcal{P}$ denote the ordered set of all patterns, where the order is the ”natural” one inherited after identifying patterns with integer numbers using the binary representation. Let ${\bf n}:\mathcal{P}\to\mathbb{N}$ be pattern counts ${\bf n}=(n_{\alpha})_{\alpha\in\mathcal{P}}$ such that all but a finite number of counts are zero, and let ${\bf n}!:=\prod_{\alpha\in\mathcal{P}}n_{\alpha}!$ and $|{\bf n}|:=\sum_{\alpha}n_{\alpha}$ the total number of patterns.
@@
$|n_{1}=2;n\rangle$ $:=$ $\frac{1}{\sqrt{2}}\left(A_{1}^{*}(n)\right)^{2}|\Omega_{n}\rangle$
@@
$|n_{11}=1;n\rangle$ $:=$ $A_{11}^{*}(n)|\Omega_{n}\rangle$
@@
5 Limit distribution of quadratures and number operators.
In this section we analyse the structure of the output state obtained by post-processing the output of a QMC with a coherent absorber, as described in section 3.1. Motivated by the fact that the statistical uncertainty scales as $1/\sqrt{n}$ we choose the absorber parameter $\theta_{0}$ to be fixed and known, and write the system parameter as $\theta=\theta_{0}+u/\sqrt{n}$, where $u$ is an unknown local parameter. For a more in depth motivation we refer to local asymptotic normality theory in appendix A.
Since the output state becomes stationary for long times, it is natural to focus on the state of the translationally invariant modes introduced in section 4. In Theorem 2 we show that asymptotically with $n$, the restricted state of these modes is a joint coherent state whose amplitude depends linearly on $u$, i.e. a Gaussian shift model. Moreover, Corollary 2 shows that the QFI of this model is equal to the QFI rate $f_{\theta_{0}}$ of the output state characterised in Theorem 1. This means that the TIMs capture the entire QFI of the output. Together with Theorem 3 of section 6, these results will be the theoretical underpinning the estimator proposed in section 7.
We consider the system and absorber together as an open system with space $\mathcal{H}_{sa}\cong\mathbb{C}^{D}$ with $D=d^{2}$, interacting with the noise units. The corresponding unitary is $W_{\theta}=V_{\theta_{0}}U_{\theta}$ where the absorber parameter is a fixed value $\theta_{0}$ (which later will be determined based on a preliminary estimation procedure) and $\theta$ is the unknown parameter to be estimated. We distinguish between the system Kraus operators $K_{\theta,i}=\langle i|U_{\theta}|0\rangle\in B(\mathcal{H}_{s})$ and the system and absorber Kraus operators $\tilde{K}_{\theta,i}=\langle i|W_{\theta}|0\rangle\in B(\mathcal{H}_{sa})$, and similarly between the system transition operator $\mathcal{T}_{\theta}$ and the system and absorber transition operator $\tilde{\mathcal{T}}_{\theta}(X)=\sum_{i=0}^{1}\tilde{K}_{\theta,i}^{*}X\tilde{K}_{\theta,i}$.
@@
    $\tilde{\mathcal{T}}_{\theta}$ has a unique invariant state $\tilde{\rho}^{\rm ss}_{\theta}>0$ and is aperiodic for $\theta$ in a neighborhood of $\theta_{0}$;
@@
Note that properties $1.$ and $2.$ in Hypothesis 2 imply that $\dot{\tilde{\rho}}^{\rm ss}=\tilde{\mathcal{R}}_{*}\dot{\tilde{\mathcal{T}}}_{*}(\tilde{\rho}^{\rm ss})$ where $\tilde{\mathcal{R}}$ is the Moore-Penrose inverse of ${\rm Id}-\tilde{\mathcal{T}}$. In addition, $\tilde{K}_{0}{|{\chi^{\rm ss}}\rangle}=\tilde{K}^{*}_{0}{|{\chi^{\rm ss}}\rangle}={|{\chi^{\rm ss}}\rangle}$. Indeed by construction $\tilde{K}_{0}|\chi^{\rm ss}\rangle=|\chi^{\rm ss}\rangle$ and $\tilde{K}_{1}|\chi^{\rm ss}\rangle=0$, and by applying $\tilde{K}_{0}^{*}\tilde{K}_{0}+\tilde{K}_{1}^{*}\tilde{K}_{1}=\mathbf{1}$ to $|\chi^{\rm ss}\rangle$ we get $\tilde{K}_{0}^{*}|\chi^{\rm ss}\rangle=|\chi^{\rm ss}\rangle$.
@@
To formulate our results below we need to introduce several superoperators acting on the system and absorber space. For $x\in B(\mathcal{H}_{sa})$, we define
[…]
@@
The upshot of this section is that (asymptotically in $n$) the statistical information of the output state is concentrated in the TIMs and the state’s restriction to the TIM Bosonic algebra is a coherent state. Formally, in order to optimally estimate the parameter, one would only need to measure the appropriate quadrature of the Gaussian shift model, as explained in section 2. However, it is not obvious how to perform such a measurement, and the theoretical insight does not seem to help on the practical side. Surprisingly, it turns out that the standard sequential counting measurement is an effective joint measurement of all the TIMs’ number operators! This will be the main result of the next section, which in conjunction with the displaced null strategy discussed in section A, provides the ingredients of a counting-based estimation strategy.
6 Limit theorem for counting trajectories
@@
Let $\gamma$ be an arbitrary but fixed real number satisfying $0<\gamma<1$. Let ${\bf m}=\{m_{\alpha}\}_{\alpha\in\mathcal{P}}$ be a set of pattern excitation counts where all occupation numbers are zero except $(m_{\alpha^{(1)}},\dots,m_{\alpha^{(k)}})$; we define $B_{\bf m}(n)$ as the set of all binary sequences of length $n$ containing $m_{\alpha^{(1)}}$ copies of $\alpha^{(1)}$s, up to $m_{\alpha^{(k)}}$ copies of $\alpha^{(k)}$s in any order, and such that between two consecutive patterns there are at least $n^{\gamma}$ 0s. We remark that $B_{\bf m}(n)$ is the empty set for every $n$ strictly smaller than $\sum_{i=1}^{k}m_{\alpha^{(i)}}|\alpha^{(i)}|+(k-1)n^{\gamma}$ ($|\alpha|$ is the length of the pattern $\alpha$).
For instance, let us consider the set of pattern excitation counts $\mathbf{m}=(2_{(11)},1_{(101)})$, given by $2$ excitations of the type $(11)$ and $1$ of type $(101)$. The set $B_{\mathbf{m}}(n)$ is empty up to $n\geq 7+2\cdot n^{\gamma}$, then it contains all the strings of length $n$ of the following form:
$0\cdots 0\mathbf{101}\underbrace{0\cdots 0}_{n_{1}}\mathbf{11}\underbrace{0\cdots 0}_{n_{2}}\mathbf{11}0\cdots 0,$
$0\cdots 0\mathbf{11}\underbrace{0\cdots 0}_{n_{1}}\mathbf{101}\underbrace{0\cdots 0}_{n_{2}}\mathbf{11}0\cdots 0,$
$0\cdots 0\mathbf{11}\underbrace{0\cdots 0}_{n_{1}}\mathbf{11}\underbrace{0\cdots 0}_{n_{2}}\mathbf{101}0\cdots 0,$
where $0\cdots 0$ stays for a sequence of all 0s and $n_{2},n_{3}\geq n^{\gamma}$.
The following Theorem shows that the distribution of the pattern counts $\bf m$} converges to a product of Poisson distributions with the same intensities as those of the number operators of the TIM in Theorem 2. This means that performing a standard output measurement and extracting the pattern counts provides an effective joint measurement of the number operators of the TIMs. This finding is essential in constructing an optimal estimator in section 7.
@@
For every positive constant $C>0$ and finite collection of excitation patterns counts ${\bf m}=(m_{\alpha^{(1)}},\dots,m_{\alpha^{(k)}})$, the following limit is equal to zero:
$\lim_{n\rightarrow\infty}\sup_{|u|<C}\left|\nu_{u,n}\left(B_{\bf m}(n)\right)-e^{-\lambda_{tot}u^{2}}\prod_{i=1}^{k}\frac{\left(\lambda_{\alpha^{(i)}}u^{2}\right)^{m_{\alpha^{(i)}}}}{m_{\alpha^{(i)}}!}\right|.$ (21)
@@
$\lambda_{tot}$ $:=$ $-\frac{1}{2}{\langle{\chi^{\rm ss}}|}(2\dot{{\cal B}}_{0*}{\cal R}_{0*}\dot{{\cal B}}_{0*}+\ddot{{\cal B}}_{0*})({|{\chi^{\rm ss}}\rangle}{\langle{\chi^{\rm ss}}|}){|{\chi^{\rm ss}}\rangle}$ (22)
@@
The previous result has some relevant consequences. Let us define the “pattern extraction” function which associates to each trajectory $\omega\in\{0,1\}^{n}$ a set of pattern counts $\{N_{\alpha,n}(\omega):\alpha\in\mathcal{P}\}\in\mathbb{N}^{\mathcal{P}}$, which is uniquely determined by the condition that $\omega$ is a maximal union of contiguous patterns separated by sequences of 0s of length at least $n^{\gamma}$ with a fixed $0<\gamma<1$; moreover, let us consider the stochastic process given by the infinite collection of independent random variables $\{N_{\alpha}:\alpha\in\mathcal{P}\}$ where $N_{\alpha}$ is a Poisson random variable with parameter $\lambda_{\alpha}u^{2}$.
@@
However, in the following we will be interested in local parameters with growing size, i.e. $|u|\leq n^{\epsilon^{\prime}}$ for some $0<\epsilon^{\prime}<1/2$. In this case we can show the following result.
@@
For $0<\epsilon^{\prime}<1/6$ and for every finite collection of excitation patterns counts $\bf m$} the following holds true:
$\lim_{n\rightarrow+\infty}\sup_{|u|\leq n^{\epsilon^{\prime}}}\left|\frac{\nu_{u,n}\left(B_{\mathbf{m}}(n)\right)}{e^{-\lambda_{tot}u^{2}}\prod_{i=1}^{k}\frac{\lambda_{\alpha^{(i)}}^{m_{i}}u^{2m_{i}}}{m_{i}!}}-1\right|=0.$ (23)
The Proof of Proposition 2 can be found in Appendix E as well. Upgrading this result to a weak convergence one similar to Corollary 3 remains the subject of future research.
7 Pattern counting estimator
In this section we describe our adaptive estimation scheme which exploits the asymptotic results presented in sections 5 and 6. The scheme involves four key ingredients: